OpenStudy (anonymous):

4 sinx/cosx - 1/cos^2x=0 How would i simplify this to: 4sinxcosx=1

OpenStudy (anonymous):

pls help

OpenStudy (anonymous):

\[\frac{ 4 \sin x }{ \cos x }-\frac{ 1 }{ \cos^2 x }=0\] is that your question?

OpenStudy (anonymous):

the 4 is seperated from sin/cos

OpenStudy (anonymous):

4(sinx/cosx)

OpenStudy (anonymous):

i typed in the question wrong sorry

OpenStudy (anonymous):

simplify it to 4(sinx)(cosx)=1

OpenStudy (dessyj1):